OpenStudy (theedwardsfamily):

What is the distance between the two points on a number line? 1. X = 4, Y = –17 2. P = 23, Q = 8 3. U = –1, V = –14 4. J = –3, K = 18 A. 13 B. 15 C. 21

OpenStudy (theedwardsfamily):

@TheShadyLP

Directrix (directrix):

This is a matching question?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Oh ok

OpenStudy (theedwardsfamily):

@Directrix yes

Directrix (directrix):

X = 4, Y = –17 Use the Ruler Postulate. The distance between two points on a number line is the absolute value of the difference of the two coordinates. On the first problem, abs of (4 - (-17) ) = abs(4 + 17) = abs(21) = 21 -->

Directrix (directrix):

Think about it this way. From -17 moving to the right to zero is 17 units. Then there are 4 more which gives 21 for the first problem.

OpenStudy (theedwardsfamily):

ok im trying to understand

Directrix (directrix):